OpenStudy (anonymous):

log_{sqrt{3}} 9 + log_{25} 5

zepdrix (zepdrix):

Hey Lisa :) Hmm interesting base on that first one :d

zepdrix (zepdrix):

\[\large\rm \log_{\sqrt3}9+\log_{25}5\]

zepdrix (zepdrix):

Is that what the problem looks like?

OpenStudy (fibonaccichick666):

This may help you http://www.purplemath.com/modules/logs.htm You can read that for a full description of what logs are and the relationship. What will be important for you is this relation: \[\huge { {y=b^x \rightarrow log_b(y)=x}}\]

OpenStudy (anonymous):

yes that is what it looks like

OpenStudy (anonymous):

sorry about the long response didn't realize how long I was away for

OpenStudy (fibonaccichick666):

so, what is your b=? y=? for \(\log_{\sqrt3}9\)?

OpenStudy (anonymous):

I'm not sure

zepdrix (zepdrix):

@lisamath12 it turns out that \(\rm 9\) is the product of a bunch of \(\rm \sqrt3\)'s. How many square root 3's would you multiply together to get 9? \(\large\rm \sqrt3\sqrt3=3\) Hmm so two of them is not enough.

OpenStudy (disqin):

\(\log_{\sqrt{3}} 9 + \log_{25} 5\) Since \(\log_{\sqrt{3}} 9 = 2\log_{3} 3^2=2\times 2\log_33=4\) Similarly \(\log_{25} 5 = 2\log_{5} 5=2\) Hence \(\log_{\sqrt{3}} 9 + \log_{25} 5=4+2=6\)

OpenStudy (unklerhaukus):

Use the change of base formula\[\boxed{\log_ba = \frac{\log_c a}{\log_c b}}\] \[\log_{\sqrt{3}} 9 + \log_{25} 5\\ = \frac{\log 9}{\log \sqrt 3}+\frac{\log5}{\log25}\\ = \frac{\log 3^2}{\log3^{1/2}}+\frac{\log5}{\log5^2}\] Then use\[\boxed{\log_b a^n=n\log_b a}\] \[= \frac{2\log 3}{\frac12\log3}+\frac{\log5}{2\log5}\] cancel the logs \[= \frac{2}{(\frac12)}+\frac12\] simplify \[=\]

OpenStudy (anonymous):

\[\log_{\sqrt{3}} 9=x,\left( \sqrt{3} \right)^x=9,\left( 3 \right)^{\frac{ x }{ 2 }}=3^2,\frac{ x }{ 2 }=2,x=4\] \[\log_{25} 5=y,25^y=5,\left( 5^2 \right)^y=5,5^{2y}=5^1,2y=1,y=\frac{ 1 }{ 2 }\] add and get the result.